# Value Residual Learning For Alleviating Attention Concentration In Transformers

### v2

@@ 4.1.3 Entropy For Analyzing Attention Concentration Effects
where $\bm{a}_{i}^{\prime}={\bm{a}_{i}}/\Big{(}{\sum_{i=1}^{l}{\bm{a}_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
@@ 4.1.4 Spectral Decomposition For Analyzing Representations
$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{i=j}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.$ (12)
where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $i=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
@@
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 5, ResFormer consistently outperforms vanilla Transformer throughout training for different models sizes. Additionally, the results in Fig. 1 (Left) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements. We further test the variant of Resformer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$. As shown in Fig.6, Resformer can accommodate a wide range of $\bm{\lambda}$ values and the performance improves as $\bm{\lambda}$ increases, achieving the best results at ${\bm{\lambda}}=2$. Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers. It suggests that the success of Resformer lies in the use of $\mathbf{V}_{1}$ and the transformation through $\mathbf{A}_{n}$. The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.     [Uncaptioned image] Figure 6: Ablation study of different $\bm{\lambda}$ for Resformer.
@@ 4.3 Ablation study of residual connection
(a) Token importance.
(b) Norms of value states.
(c) Norms of hidden states.
Figure 10: The token importance, value-state norms, and hidden-state norms of the first token in each sequence across layers in ResFormer vs. the vanilla Transformer.
(a) Transformer.
(b) NeuTRENO.
(c) Resformer.
In Eqn. 4, we employ residual connections for the values. We compare this approach with models that add residual connections to queries or keys. The results, shown in Fig. 9, indicate that only residual connections for values yield positive effects. One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys can significantly impact attention matrix.
Moreover, we compare with the models based on Eqn. 2 and Eqn. 6. From the cross layer attention perspective, the former uses a gold attention matrix in Eqn. 2 to map historical values, while the latter uses an all-ones matrix in Eqn. 6. The results in Fig. 9 align with Fig. 4, showing that identity mapping causes significant perturbations, leading to poor performance. Interestingly, using current attention as the mapping matrix results in an even lower final loss than using gold attention.
@@ 4.4 Downstream Evaluations
We compare the different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way. The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020). The results in Table 1 show that ResFormer (82M) achieves an average accuracy improvement of nearly 3% compared to the vanilla Transformer (82M).
Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg.
@@
Figure 12: Left: Distribution of eigenvalues for the value states in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
@@ 4.5 Visualization of ResFormer
Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data. Fig. 1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer. We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2. Given that attention clustering often occurs on the first token, we primarily show its results in Fig. 10. The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2023), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024). Guo et al. (2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains. Specifically, for tokens lacking semantic information, such as the first token, a large value state magnitude can adversely affect the prediction of future tokens if it is overly attended to, resulting in higher loss.
@@
Results in Fig. 14 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is much faster than vanilla attention. However, as the training sequence length increases, the SVFormer model performs better. In this way, we focus on the critical point, defined as the number of training steps exceeded. Fig. 14 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend. We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
@@ 4.7 Other Factors Influencing SVFormer
Except for the 2M model, Llama-like models ranging from 82M to 468M, as well as models with the GPT2 architecture, exhibit similar critical points and final losses (see Fig. 15(c) and Fig. 15(d)). This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to model size and architecture.
@@
(c) Model Size.
(d) Architecture.
Figure 15: The relative training loss for SVFormer under different hyper-parameter setting.
@@ 5 Conclusion
In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers. From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value states of the current layer and those of the first layer before the attention operation to alleviate attention concentration. Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half. We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
Acknowledgement
This work was supported by the Research Center for Industries of the Future at Westlake University (Grant No. WU2023C017) and the Key Research.
@@ Acknowledgement
We have detailed the complete experiments setup such as batch size, optimizer and learning rates in Section 4.1.1. Besides, we have released the source codes. These resources should be sufficient to reproduce the results of the paper.
@@ A.1 Token Similarity Analysis
where $\{\bm{h}_{i}\}_{i=1}^{l}$ is the hidden state of the $i$-th token and $\operatorname{Sim}(\cdot)$ denotes the operation of cosine similarity. The results in Fig. 18 are align with the results in Fig. 1. In the case of Llama and Mistral, the average token similarity demonstrates an “M”-shaped pattern with increasing network depth, while entropy follows a “W”-shaped pattern at corresponding positions. These trends suggest that attention concentration indeed leads to the phenomenon of over-smoothing.
@@ A.2 Attention Concentration Visualization
We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers. The results are averaged from 1,000 different sequences so that only the start token is the same and special across all sequences. Fig. 19 (First row) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO. For Resformer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention. Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers. In contrast, for Resformer, the value state norm of the start token is approximately 0.7 times that of other tokens, while the peak hidden state norm is about three times larger.
Moreover, We conduct an entropy analysis on test examples of varying lengths. The results in Fig.20 indicate that Resformer is more effective in mitigating attention concentration when the sequences are shorter.
Token
importance
Norms of
value states
Norms of
hidden states
@@
Figure 19: Visualization of the token importance, norms of value states and norms of hidden states.
(a) Length: 20 tokens.
(b) Length: 100 tokens.
(c) Length: 500 tokens.
(d) Length: 2,048 tokens.
Figure 20: The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer for sequence with different lengths.

### v3

@@ 4.1.3 Entropy For Analyzing Attention Concentration Effects
where $a_{i}^{{\prime}_{i}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
@@ 4.1.4 Spectral Decomposition For Analyzing Representations
$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{j=1}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.\vspace{-3mm}$ (12)
where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $j=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
Figure 6: (Left) Validation loss as model size scales from 82M to 468M parameters. (Right) Validation loss for the 468M parameter model evaluated every 2B tokens. ResFormer achieves approximately 10.4%-13.6% reduction in both model parameters and training data.
Figure 7: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths. (Right) Relative training loss of various Transformer variants compared to the vanilla Transformer model, with model size fixed at 82M parameters.
@@
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 7 (Left), ResFormer consistently outperforms vanilla Transformer throughout training across different training sequence lengths. Additionally, the results in Fig. 7 (Right) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
We also analyzed how ResFormer and Transformer scale at model size and data size. ResFormer and Transformer are trained on similar experiment setting. On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set. As shown in Fig.7 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 10.4% fewer model parameters. On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 13.6% fewer training tokens to achieve the same validation loss as Transformer. The validation loss for these models is available in Appendix A.6.
We further test the variant of ResFormer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$. As shown in Fig.8, ResFormer can accommodate a wide range of $\bm{\lambda}$ values and the performance improves as $\bm{\lambda}$ increases, achieving the best results at ${\bm{\lambda}}=2$. Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers. It suggests that the success of ResFormer lies in the use of $\mathbf{V}_{1}$ and the mapping by $\mathbf{A}_{n}$. The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.
Figure 8: Ablation study of different $\bm{\lambda}$ for ResFormer.
@@ 4.3 Ablation study of residual connection
In Eqn. 4, we employ residual connections for the values. We compare this approach with models that add residual connections to queries or keys. The results, shown in Fig. 11, indicate that only residual connections for values yield positive effects. One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys significantly impacts it.
Moreover, we compare with the models based on Eqn. 2 and Eqn. 6. The results in Fig. 11 align with Fig. 4, showing that identity mapping causes significant perturbations, leading to poor performance. Interestingly, ResFormer achieves an even lower final loss than ResFormer. It suggests that ResFormer’s impact on the attention optimization is better by mitigating value-state drains.
@@ 4.4 Downstream Evaluations
We compare the different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way. The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020). The results in Table 1 show that ResFormer achieved an average accuracy improvement of nearly 3% compared to the vanilla Transformer.
Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg
@@
Figure 12: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
@@ 4.5 Visualization of ResFormer
We also analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data. Fig. 1 (Right) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer. We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2. Given that attention clustering often occurs on the first token, we primarily show its results in Fig. 2. The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2024), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024). Guo et al. (2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains. Specifically, for tokens lacking semantic information like start tokens, a large value state magnitude can adversely affect the prediction of subsequent tokens if they are overly attended to. When there is no value-state drains, models will reduce attention clustering to these tokens to minimize loss.
@@
Results in Fig. 14 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is faster than vanilla Transformer at the early stage. However, as the training sequence length increases, the SVFormer model performs better. In this way, we focus on the critical point, defined as the number of training steps exceeded. Fig. 14 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend. We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
@@ 4.7 Other Factors Influencing SVFormer
Llama-like models and GPT2-like models exhibit similar critical points and final losses (see Fig. 15(c)). This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to architecture. Compared with Transformer, SVFormer requires a 12.2% increase in parameters to achieve the same validation loss while reducing the $KV$-cache by nearly half.
@@
(c) Architecture.
(d) Model Size.
Figure 15: The relative training loss for SVFormer under different hyper-parameter setting and the validation loss as model size scales from 82M to 468M parameters.
@@ 5 Conclusion
In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers. From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration. Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half. We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
@@
We have detailed the complete experiments setup such as batch size, optimizer, learning rates in Section 4.1.1. Besides, we will release source codes once our paper is made public. These resources should be sufficient to reproduce results of the paper.
@@ A.1 Token Similarity Analysis
where $\{\bm{h}_{i}\}_{i=1}^{l}$ is the hidden state of the $i$-th token and $\operatorname{Sim}(\cdot)$ denotes the operation of cosine similarity. The results in Fig. 18 are align with the results in Fig. 1. In the case of Llama and Mistral, the average token similarity demonstrates an “M”-shaped pattern with increasing network depth, while entropy follows a “W”-shaped pattern at corresponding positions. These trends indicate a strong correlation between attention concentration and over-smoothing.
@@ A.2 Attention Concentration Visualization
Transformer
NeuTRENO
ResFormer
SVFormer
(a) Token importance.
(b) Value-state norms.
(c) Hidden-state drains.
Figure 19: Visualization of token importance, value state norms, and hidden state norms across different token positions and layers in 468M models.
(a) Entropy.
(b) Token importance.
(c) Value-state norms.
(d) Hidden-state norms.
Figure 20: Token analysis in 82M models. (a) Importance entropy of sequences; (b) Token importance, (c) value-state norms, and (d) hidden-state norms of the first token across layers.
@@
(d) SVFormer.
We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers. The results are averaged from 1,000 different sequences so that only the start token is the same and special across all sequences. Fig. 19 (First column) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO. For ResFormer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention. Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers. In contrast, for ResFormer, the value state norm of the start token exceeds half the magnitude of other tokens, while the peak hidden state norm is less than triple the average. Fig.21 further illustrates the distribution of token importance, where TOP-$i$ represents the $i$-th largest token importance within a sequence. Compared to Transformer and NeuTRENO, ResFormer and SVFormer exhibit a more uniform distribution of token importance.
Similar to Fig.2, we conducted experiments on 82M models, with results shown in Fig.20. We also illustrate the attention pattern of cross-layer attention introduced in Eqn. 2. The results demonstrate that while cross-layer attention successfully mitigates the problem of attention concentration, it still exhibits value-state drains.